OpenStudy (volleyballlover55):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 9 inches. The height of the cone is 18 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (volleyballlover55):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Volume of a Cone \[\Large V = \frac{1}{3}\pi*r^2*h\] --------------------------------------- Volume of Cylinder \[\Large V = \pi*r^2*h\] ---------------------------------------- V = volume pi = 3.14 approximately r = radius h = height

OpenStudy (volleyballlover55):

ok...

jimthompson5910 (jim_thompson5910):

do those formulas look familiar?

OpenStudy (volleyballlover55):

yes..

jimthompson5910 (jim_thompson5910):

ok tell me what you get for the volume of the cylinder

OpenStudy (volleyballlover55):

ok...

OpenStudy (volleyballlover55):

whats the radius and height?

jimthompson5910 (jim_thompson5910):

`A cylinder and a cone have the same diameter: 8 inches.` diameter = 8 radius = 4 (cut the diameter in half to get the radius)

OpenStudy (volleyballlover55):

ok!

jimthompson5910 (jim_thompson5910):

`The height of the cylinder is 9 inches`

OpenStudy (volleyballlover55):

i got v=452.16

jimthompson5910 (jim_thompson5910):

`i got v=452.16` me too

jimthompson5910 (jim_thompson5910):

now find the volume of the cone

OpenStudy (volleyballlover55):

ok... but for the steps on how i found that answer..

jimthompson5910 (jim_thompson5910):

what about them specifically?

OpenStudy (volleyballlover55):

well can you tell me the steps you got to find the answer of v=452.16

jimthompson5910 (jim_thompson5910):

you tell me. Give it a shot and post what you think

OpenStudy (volleyballlover55):

ok ill try!

OpenStudy (volleyballlover55):

sorry, give me a sec.. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

take your time

OpenStudy (volleyballlover55):

thx :)

OpenStudy (volleyballlover55):

V=3.14×16×9 V=452.19

OpenStudy (volleyballlover55):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

good, before that I would have these steps V = pi*r^2*h V = 3.14*4^2*9

jimthompson5910 (jim_thompson5910):

so all together, your steps would look something like this V = pi*r^2*h V = 3.14*4^2*9 V=3.14*16*9 V=452.19

jimthompson5910 (jim_thompson5910):

btw I think you mean to say 452.16

jimthompson5910 (jim_thompson5910):

so all together, your steps would look something like this V = pi*r^2*h V = 3.14*4^2*9 V=3.14*16*9 V=452.16

OpenStudy (volleyballlover55):

oh yes i did sorry.. perfect! is this the complete answer for that question :)

OpenStudy (volleyballlover55):

V = pi*r^2*h V = 3.14*4^2*9 V=3.14*16*9 V=452.16

OpenStudy (volleyballlover55):

?

jimthompson5910 (jim_thompson5910):

That only takes care of the cylinder. Now you have to do the cone.

OpenStudy (volleyballlover55):

ok..

OpenStudy (volleyballlover55):

whats the radius and height of the cone?

jimthompson5910 (jim_thompson5910):

it's given in the instructions

OpenStudy (volleyballlover55):

ok!

OpenStudy (volleyballlover55):

4 and 18?

jimthompson5910 (jim_thompson5910):

correct on both

OpenStudy (volleyballlover55):

13×3.14×16×18=301.44

jimthompson5910 (jim_thompson5910):

it's 1/3 and not 13

OpenStudy (volleyballlover55):

sorry i forgot to put the 1/3

jimthompson5910 (jim_thompson5910):

volume of cone V = (1/3)*pi*r^2*h V = (1/3)*3.14*4^2*18 V = (1/3)*3.14*16*18 V = 301.44

OpenStudy (volleyballlover55):

ok awesome! now what?

jimthompson5910 (jim_thompson5910):

volume of cylinder V = 452.16 volume of cone V = 301.44 Divide the two values (big over small)

OpenStudy (volleyballlover55):

ok!

OpenStudy (volleyballlover55):

i got 1.5

jimthompson5910 (jim_thompson5910):

me too

OpenStudy (volleyballlover55):

ok so ....

jimthompson5910 (jim_thompson5910):

`What is the relationship between the volume of this cylinder and this cone?` the volume of the cylinder is 1.5 times bigger than the volume of the cone

OpenStudy (volleyballlover55):

ok perfect!

OpenStudy (volleyballlover55):

so the complete answer for that question is ::: The volume of the cylinder is: V = pi*r^2*h V = 3.14*4^2*9 V=3.14*16*9 V=452.16 The Volume of the cone is: V = (1/3)*pi*r^2*h V = (1/3)*3.14*4^2*18 V = (1/3)*3.14*16*18 V = 301.44 452.16 / 301.44=1.5 The volume of the cylinder is 1.5 times bigger than the volume of the cone.

OpenStudy (volleyballlover55):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I would add in one thing

jimthompson5910 (jim_thompson5910):

The volume of the cylinder is: V = pi*r^2*h V = 3.14*4^2*9 V=3.14*16*9 V=452.16 The Volume of the cone is: V = (1/3)*pi*r^2*h V = (1/3)*3.14*4^2*18 V = (1/3)*3.14*16*18 V = 301.44 (volume of cylinder)/(volume of cone) = 452.16 / 301.44=1.5 The volume of the cylinder is 1.5 times bigger than the volume of the cone.

OpenStudy (volleyballlover55):

ok perfect!!!

OpenStudy (volleyballlover55):

omg! yay we're done :) your the best you know!

jimthompson5910 (jim_thompson5910):

glad to be of help